# Assessing differences between local galaxy dust attenuation and point source extinction within the same environments

### v1

@@ 2.2 Spiral Galaxy Model
The light source system for the spiral galaxy is divided into two components: a disk and a bulge. The disk is described by a disk geometry characterized by a double-exponential profile given by Eq. 3:
@@
The bulge component is described by a flattened spheroidal geometry, defined by Eq. 1 with a flattening parameter of $q=0.7$ and a Sérsic (1963) $\rho(r)$ density profile given by Eq. 4:
@@
The dust medium geometry is taken to be the same as that of the spiral source disk, described by a disk characterized by a double-exponential profile as defined by Eq. 3, with the normalization $\rho_{0}$ being varied to produce simulated galaxies with different optical depths. We use a scale length parameter $h_{R}=4000$ pc and a scale height parameter $h_{z}=350$ pc. Even though the dust disk is typically found to be thinner than the stellar disk (e.g., De Geyter et al., 2014), we opt for using disks with the same thickness. We do so because we are primarily interested in the interaction between light and dust and thus want to limit the number of stars effectively outside the dust disk. However, we find that using a thinner dust disk does not significantly alter the results discussed below.
@@ 2.3 Elliptical Galaxy Model
Once more following Baes et al. (2000), the dust system for the elliptical galaxy is described by a flattened spheroidal geometry, defined by Eq. 1 with a flattening parameter of $q=0.5$ and a Plummer (1911) $\rho(r)$ density profile, as described by Eq. 2 with $c=4000$ pc. The normalization $\rho_{0}$ is varied to produce different optical depths in the dust system.
@@ 3.1 Point Source Model
In this section we present an analysis of simulated photometric observations for the point source model. Particularly, we analyze the impact of different apertures on the dust properties recovered for the model and the transition between extinction and attenuation regimes.
The observed flux can be divided into two components: 1) the direct flux, comprising photon packets that reach the detector without interacting with dust; 2) the scattered flux, comprising packets that have gone through at least one scattering event, ending up in the line of sight. Fig. 1 shows the ratio between these two fluxes for face-on observations of a point source inside a galactic dust cloud as seen by simulated instruments with different aperture radii. The dust density profile was normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$.
Figure 1: Ratio between the scattered (into the line of sight) and direct fluxes as a function of the wavelength $\lambda$ for face-on observations of a simulated point source, immersed in a galactic dust cloud. Simulated observations for a variety of aperture radii are shown. The cloud dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$.
@@
This is further illustrated in Fig. 2, which shows, as an example, the values of $A_{V_{eff}}$ and $R_{V_{eff}}$ for a variety of observation angles when using two different simulated instruments, with aperture radii of $1$ AU (left) and $20$ kpc (right), respectively. As hinted before, the observed behavior greatly depends on the aperture used. For the smaller aperture, we recover a constant value of $R_{V}=3.068$ that does not depend on the observation angle and matches exactly the value of the extinction curve slope for the Zubko dust mix $R_{V_{Zubko}}$, as directly computed from the material extinction cross-sections outputted by SKIRT. Likewise, the values of $A_{V}$ can be taken as true indicators of dust extinction, as they correctly reflect the amount of light that has been absorbed or scattered away from the line of sight.
@@
For the larger aperture, we note that the recovered $A_{V_{eff}}$ values are $57\%$ to $43\%$ lower than the respective true extinction $A_{V}$. This is a result of the scattered light that is now contributing to the observations, as seen if Fig. 1. Moreover, we note that the relation between $A_{V}$ and $A_{V_{eff}}$ is close to linear, indicating that within the same dust cloud the scattered and absorbed fluxes scale consistently at this wavelength. Additionally, we find that $R_{V_{eff}}$ is highly dependent on the galaxy inclination, ranging from $R_{V_{eff}}=2.212$, for face-on observations, to $R_{V_{eff}}=3.120$, for edge-on observations. We conclude that, for this particular example, an increase in the observation aperture can result in a $28\%$ deviation from the intrinsic optical extinction $R_{V}$.
@@
Figure 2: Values of effective $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ for a simulated point source immersed in a galactic dust cloud. Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). The results from two different instruments are presented: One with an aperture radius of $1$ AU (left) and another with an aperture radius of $20$ kpc (right). The cloud dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$.
@@
Figure 3: Values of effective $R_{V_{eff}}$ as a function of aperture radius for a simulated point source immersed in a galactic dust cloud. Results are shown in blue for $\tau_{V_{X}}=5$, orange for $\tau_{V_{X}}=15$ and green for $\tau_{V_{X}}=25$. Face-on results are represented by dots, while edge-on results are represented by squares. The dashed black line reflects the true $R_{V}$ of the medium.
@@
We can refine our analysis by looking at the $R_{V_{eff}}$ values recovered for each of the apertures discussed in Fig. 1, which are plotted in Fig. 3. In addition to the $\tau_{V_{X}}=5$ normalization discussed above, results for two other dust densities, $\tau_{V_{X}}=15$ and $\tau_{V_{X}}=25$, are presented, for both face-on and edge-on observations. We find that, for the current setup, any aperture with a radius larger than $25-50$ pc results in a $R_{V_{eff}}$ that deviates from $R_{V_{Zubko}}$. In particular, as stated above, we confirm that galaxy inclination has a great impact on both $A_{V_{eff}}$ and $R_{V_{eff}}$, as face-on and edge-on simulated observations differ greatly from each other.
@@
The differences in the simulated $R_{V_{eff}}$ arise from the interplay between the direct and scattered fluxes. In general, the scattered flux tends to dominate the bluer wavelengths, with the direct flux dominating the redder wavelengths, as shown in Fig. 4. As we move from a face-on view to an edge-on view, the scattered flux becomes even bluer when compared to the direct flux. Overall, this results in a flatter edge-on attenuation curve, boasting higher values of $R_{V_{eff}}$ than those observed for the face-on view, as seen in Fig. 3. This effect is primarily driven by the thickening of the optical light path for the edge-on observations, which increases the interaction probability between light and the medium, contributing to the absorption and, most relevant, the scattering of more blue light away from the direct flux. Although this increase in optical depth is also expected to make the scattered flux overall redder, this component still becomes bluer when compared to the direct flux. Likewise, if the dust density is increased, we expect an increase in $R_{V_{eff}}$ across all observation angles, which is precisely what is observed from the simulations. As such, a higher dust column density actually brings the face-on $R_{V_{eff}}$ values closer to $R_{V_{Zubko}}$, while it increases the deviation for the edge-on values.
@@
Figure 4: Flux components for face-on (top) and edge-on (bottom) simulated observations of a point source immersed in a galactic dust cloud with $\tau_{V_{X}}=5$. The observations were simulated for an aperture radius of $r=20$ kpc. The emitted flux, total observed scattered and direct fluxes are shown in purple, blue, orange and green, respectively.
@@
We also note that our point source model is not representative of SN observations. While the angular resolution of most ground based SN Ia surveys would place them above our threshold for the extinction regime, the fact is that, due to their short lifetimes, light can only effectively travel and scatter within a small vicinity around the transient (roughly $0.1$ pc in the first four months after explosion). This means that, unless surrounded by a very dense and exotic dust cloud that could give rise to light echoes, as described in Wang et al. (2008) and Bulla et al. (2018), the SN-dust interaction should always be properly described by extinction. These SN light-echoes are usually only visible at very late stages in the SN evolution. However, the scattering effect described above might be important for other long-lived point sources like stars (e.g. Cepheids) or even compact stellar clusters, taking into account that their source luminosity is fainter than a SN.
@@ 3.2.1 Global Observations
Fig. 5 shows the simulated values of $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ for both a spiral and an elliptical-like galaxy model, observed from a variety of angles with an aperture radius of $20$ kpc. In both models, the dust density profile was normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$.
Figure 5: Values of $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ (top) and the corresponding values $A_{V_{dir}}$ and $R_{V_{dir}}$, calculated from only the direct flux component, ignoring the scattered flux component (bottom). Results for a simulated spiral (left) and elliptical galaxy (right) are shown. Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). An aperture radius of $20$ kpc was used. The galaxies dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$. The dashed black line reflects the true extinction $R_{V_{Zubko}}$.
@@
As mentioned in Section 3.1, the trend between $A_{V_{eff}}$ and $R_{V_{eff}}$ is primarily driven by the scattering of light into the line of sight. This scattered component, when compared to the direct flux, is redder for the face-on observations, lowering the observed $R_{V_{eff}}$, and bluer for the edge-on view, resulting in higher $R_{V_{eff}}$. In addition to this, the fact that the sources now extend to different depths and contribute to the same optical path means that $R_{V_{eff}}$ deviates even more from $R_{V_{Zubko}}$.
@@
From the above, we conclude that the reddening caused by changes in the direct optical paths behaves similarly to the reddening caused by excess scattered light, with $R_{V_{eff}}$ increasing with the observation angle. In this case, this increase is due to the fact that, as we move from a face-on to an edge-on orientation, there is an increase in the number of different optical paths in our observation, as we are probing much deeper into the dust cloud. This also explains why the spiral galaxy face-on direct flux observations in Fig. 5 produce the closest results to $R_{V_{Zubko}}$, as this is the scenario in which the sources are most concentrated, given the narrower depth of the galaxy. We also find that, as observed for the point source, the $A_{V_{eff}}$ values obtained for only the direct flux are systematically higher than those obtained from the full observed flux, as expected.
@@
This dust density effect is, however, still secondary to the variation of the observation angle, which produces a much larger change in $R_{V_{eff}}$. As such, the effective dust properties of a galaxy appear to be more a function of geometry and morphology than of the dust abundance itself. In practice, this means that, when using global galaxy aperture photometry, the true optical properties of the dust medium, namely $R_{V_{Zubko}}$, become obfuscated and cannot be easily recovered.
Figure 6: Values of $R_{V_{eff}}$ as a function of the total optical depth normalization along the x-axis $\tau_{V_{X}}$, for a spiral (left) and elliptical galaxy (right). Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). An aperture radius of $20$ kpc was used. The dashed black line reflects the true extinction $R_{V_{Zubko}}$.
In Fig. 7, we present a comparison between the $A_{V_{eff}}$-$R_{V_{eff}}$ trend obtained from the various simulated observations detailed in Fig. 6 and the trend obtained for Dark Energy Survey (DES; Abbott et al., 2018) global galaxy fits, as discussed in Duarte et al. (2023). The plot covers different galaxy types, dust densities, and orientations. We find that, with the exception of the absolute $A_{V_{eff}}$ values, our simulations are able to fully reproduce the observed behavior. The mismatch in attenuation values might actually stem from a bias in the DES fits of Duarte et al. (2023), possibly stemming from the Prospector Python package (Johnson et al., 2021). In fact, Zahid et al. (2013) and Salim et al. (2018) recover $A_{V_{eff}}$ value ranges for observed galaxies that are similar to the one obtained from our simulations. The increased level of scatter exhibited by the real data is most likely due to the fact that a real sample is expected to reflect a wide mixture of different morphologies and dust types.
Our results are also consistent with simulated observations of galaxies with different inclinations performed by Chevallard et al. (2013), as well as with other global galaxy observations (e.g Salmon et al., 2016; Leja et al., 2017; Narayanan et al., 2018). Overall, we feel confident in asserting that the observed $A_{V_{eff}}$-$R_{V_{eff}}$ relation is indeed driven by the observation angle, with factors such as dust density and galaxy type contributing mainly to an increase in the trend dispersion.
@@ 3.2.2 Local Observations
Contrary to what we found for a point source, observing with a $25$ pc resolution appears to be insufficient to meaningfully mitigate the scattered flux component across the entire galaxy, which now comes from integrated light. Even so, it appears that, for regions such as the centers of the galaxies, where the dust density is high, the scattered light is indeed negligible. In fact, looking at Fig. 9, where maps of $A_{V_{eff}}$ and $R_{V_{eff}}$ for these two galaxies are shown, we see that, for these central regions, we recover a value of $R_{V_{eff}}$ close to $R_{V_{Zubko}}$. As we move away from the center of the galaxy, we find a decrease in the values of $R_{V_{eff}}$. This is consistent with the gradient recovered by Hutton et al. (2015), who find that the observed $R_{V_{eff}}$ values decrease with the galactocentric distance, when using both ultraviolet and optical photometry. While the authors interpret this as evidence for a radial variation of intrinsic dust properties in galaxies, Fig. 9 shows that the same behavior can be recovered with a uniform dust type as a consequence of different levels of scattering and dust-to-stars distributions. There are also some outlier pixels with uncharacteristically high $R_{V_{eff}}$ values, but these arise mainly due to stochastic variations in the simulation and are particularly relevant for thin dust layers, such as the one present in the spiral edge-on view.
@@
Figure 9: Spatial maps of the values of $A_{V_{eff}}$ and $R_{V_{eff}}$, for a simulated spiral (left) and elliptical face-on galaxy (right). A pixel size corresponding to a $25$ pc resolution was used. The galaxies dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$. Only pixels with $A_{V_{eff}}>0$ and $R_{V_{eff}}>0$ are shown.
@@
Figure 10: 2D maps of the values for $A_{V_{eff}}$ and $R_{V_{eff}}$ for a face-on simulated spiral (left) and elliptical galaxy (right). The galaxies dust density has been normalized so that the total optical depth along the x-axis is $\tau_{V_{X}}=5$. Each point corresponds to a single pixel with a resolution of $25$ pc, color-coded according to their galactocentric distance. Binned medians are shown in red, with error-bars defined by the bin standard deviation. The dashed black line reflects the true $R_{V}$ of the medium, with the dotted black lines defining a 10% deviation from this value. Only pixels with $A_{V_{eff}}>0$ and $R_{V_{eff}}>0$ are shown.
@@
The relation between the two dust parameters becomes even more apparent if we look at the $A_{V_{eff}}-R_{V_{eff}}$ trends for both face-on galaxy models, plotted in Fig. 10. As stated above, in addition to the light scattered into the line of sight, which tends to be reddened and contributes to drive $R_{V_{eff}}$ down, one must also consider the existence of multiple sources at different depths in the same optical path, which, when taken together, contribute to an increase in $R_{V_{eff}}$. For higher values of $A_{V_{eff}}$, and thus higher dust densities, the amount of light that can escape from the innermost parts of the galaxy, be it scattered or otherwise, drastically goes down. As such, the observed flux becomes largely dominated by the direct flux of stars located in the outermost layers of the galaxy, which mitigates both of the above-mentioned effects.
@@
For this reason, we can recover a value of $R_{V_{eff}}$ within $10\%$ of $R_{V_{Zubko}}$ when observing areas with large optical depth near the center of the galaxies. This recovery is also aided by the high central source density, which also contributes to make the scattered component subdominant in this particular region. We end up slightly overestimating the true extinction value mainly because it is impossible to completely eliminate the contributions from sources deeper in the galaxy, which slightly increased the values of $R_{V_{eff}}$, as seen for the stellar population modeled here.
@@
Figure 11: Similar to Fig. 10 for simulated edge-on spiral (left) and elliptical galaxy (right).

### v2

@@ 2.2 Spiral galaxy model
The light source system for the spiral galaxy was divided into two components: a disk and a bulge. The disk was described by a disk geometry characterized by a double-exponential profile given by Eq. 3:
@@
The bulge component was described by a flattened spheroidal geometry, defined by Eq. 1 with a flattening parameter of $q=0.7$ and a Sérsic (1963) $\rho(r)$ density profile given by Eq. 4:
@@
The dust medium geometry was taken to be the same as that of the spiral source disk, described by a disk characterized by a double-exponential profile as defined by Eq. 3, with the normalization, $\rho_{0}$, being varied to produce simulated galaxies with different optical depths. We used a scale length parameter $h_{R}=4000$ pc and a scale height parameter $h_{z}=350$ pc. Even though the dust disk is typically found to be thinner than the stellar disk (e.g., De Geyter et al., 2014), we opted to use disks with the same thickness. We did so because we were primarily interested in the interaction between light and dust and thus wanted to limit the number of stars effectively outside the dust disk. However, we find that using a thinner dust disk does not significantly alter the results discussed below.
@@ 2.3 Elliptical galaxy model
Once more following Baes et al. (2000), the dust system for the elliptical galaxy was described by a flattened spheroidal geometry, defined by Eq. 1 with a flattening parameter of $q=0.5$ and a Plummer (1911) $\rho(r)$ density profile, as described by Eq. 2 with $c=4000$ pc. The normalization, $\rho_{0}$, was varied to produce different optical depths in the dust system.
@@ 3.1 Point source model
In this section we present an analysis of simulated photometric observations for the point source model. In particular, we analyze the impact of different apertures on the dust properties recovered for the model and the transition between extinction and attenuation regimes.
The observed flux can be divided into two components: 1) the direct flux, comprising photon packets that reach the detector without interacting with dust; 2) the scattered flux, comprising packets that have gone through at least one scattering event, ending up in the line of sight. Fig. 1 shows the ratio between these two fluxes for face-on observations of a point source inside a galactic dust cloud as seen by simulated instruments with different aperture radii. The dust density profile was normalized so that the total optical depth along the total R axis is $\tau_{V_{R}}=5$.
Figure 1: Ratio between the scattered (into the line of sight) and direct fluxes as a function of the wavelength $\lambda$ for face-on observations of a simulated point source, immersed in a galactic dust cloud. Simulated observations for a variety of aperture radii are shown. The cloud dust density has been normalized so that the total optical depth along the R axis is $\tau_{V_{R}}=5$.
@@
This is further illustrated in Fig. 2, which shows, as an example, the values of $A_{V_{eff}}$ and $R_{V_{eff}}$ for a variety of observation angles when using two different simulated instruments, with aperture radii of $1$ AU (left) and $20$ kpc (right), respectively. As hinted before, the observed behavior greatly depends on the aperture used. For the smaller aperture, we recover a constant value of $R_{V}=3.068$ that does not depend on the observation angle and matches exactly the value of the extinction curve slope for the Zubko dust mix $R_{V_{Zubko}}$, as directly computed from the material extinction cross sections outputted by SKIRT. Likewise, the values of $A_{V}$ can be taken as true indicators of dust extinction, as they correctly reflect the amount of light that has been absorbed or scattered away from the line of sight.
@@
For the larger aperture, we note that the recovered $A_{V_{eff}}$ values are $57\%$ to $43\%$ lower than the respective true extinction $A_{V}$. This is a result of the scattered light that is now contributing to the observations, as is seen in Fig. 1. Moreover, we note that the relation between $A_{V}$ and $A_{V_{eff}}$ is close to linear, indicating that within the same dust cloud the scattered and absorbed fluxes scale consistently at this wavelength. Additionally, we find that $R_{V_{eff}}$ is highly dependent on the galaxy inclination, ranging from $R_{V_{eff}}=2.212$, for face-on observations, to $R_{V_{eff}}=3.120$, for edge-on observations. We conclude that, for this particular example, an increase in the observation aperture can result in a $28\%$ deviation from the intrinsic optical extinction $R_{V}$.
@@
Figure 2: Values of effective $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ for a simulated point source immersed in a galactic dust cloud. Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). The results from two different instruments are presented: One with an aperture radius of $1$ AU (left) and another with an aperture radius of $20$ kpc (right). The cloud dust density has been normalized so that the total optical depth along the R axis is $\tau_{V_{R}}=5$.
@@
Figure 3: Values of effective $R_{V_{eff}}$ as a function of aperture radius for a simulated point source immersed in a galactic dust cloud. Results are shown in blue for $\tau_{V_{R}}=5$, orange for $\tau_{V_{R}}=15$ and green for $\tau_{V_{R}}=25$. Face-on results are represented by dots, while edge-on results are represented by squares. The dashed black line reflects the true $R_{V}$ of the medium.
@@
We can refine our analysis by looking at the $R_{V_{eff}}$ values recovered for each of the apertures discussed in Fig. 1, which are plotted in Fig. 3. In addition to the $\tau_{V_{R}}=5$ normalization discussed above, results for two other dust densities, $\tau_{V_{R}}=15$ and $\tau_{V_{R}}=25$, are presented, for both face-on and edge-on observations. We find that, for the current setup, any aperture with a radius larger than $25-50$ pc results in a $R_{V_{eff}}$ that deviates from $R_{V_{Zubko}}$. In particular, as was stated above, we confirm that galaxy inclination has a great impact on both $A_{V_{eff}}$ and $R_{V_{eff}}$, as face-on and edge-on simulated observations differ greatly from each other.
@@
The differences in the simulated $R_{V_{eff}}$ arise from the interplay between the direct and scattered fluxes. In general, the scattered flux tends to dominate the bluer wavelengths, with the direct flux dominating the redder wavelengths, as is shown in Fig. 4. As we move from a face-on view to an edge-on view, the scattered flux becomes even bluer when compared to the direct flux. Overall, this results in a flatter edge-on attenuation curve, boasting higher values of $R_{V_{eff}}$ than those observed for the face-on view, as is seen in Fig. 3. This effect is primarily driven by the thickening of the optical light path for the edge-on observations, which increases the interaction probability between light and the medium, contributing to the absorption and, most relevantly, the scattering of more blue light away from the direct flux. Although this increase in optical depth is also expected to make the scattered flux overall redder, this component still becomes bluer when compared to the direct flux. Likewise, if the dust density is increased, we expect an increase in $R_{V_{eff}}$ across all observation angles, which is precisely what is observed from the simulations. As such, a higher dust column density actually brings the face-on $R_{V_{eff}}$ values closer to $R_{V_{Zubko}}$, while it increases the deviation for the edge-on values.
@@
Figure 4: Flux components for face-on (top) and edge-on (bottom) simulated observations of a point source immersed in a galactic dust cloud with $\tau_{V_{R}}=5$. The observations were simulated for an aperture radius of $r=20$ kpc. The emitted, total observed, scattered, and direct fluxes are shown in purple, blue, orange, and green, respectively.
@@
We also note that our point source model is not representative of SN observations. While the angular resolution of most ground based SN Ia surveys would place them above our threshold for the extinction regime, the fact is that, due to their short lifetimes, light can only effectively travel and scatter within a small vicinity around the transient (roughly $0.1$ pc in the first four months after explosion). This means that, unless surrounded by a very dense and exotic dust cloud that could give rise to light echoes, as was described in Wang et al. (2008) and Bulla et al. (2018), the SN-dust interaction should always be properly described by extinction. These SN light echoes are usually only visible at very late stages in the SN evolution. However, the scattering effect described above might be important for other long-lived point sources like stars (e.g., Cepheids) or even compact stellar clusters, taking into account that their source luminosity is fainter than a SN.
@@ 3.2.1 Global observations
Fig. 5 shows the simulated values of $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ for both a spiral and an elliptical-like galaxy model, observed from a variety of angles with an aperture radius of $20$ kpc. In both models, the dust density profile was normalized so that the total optical depth along the R axis is $\tau_{V_{R}}=5$.
Figure 5: Values of $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ (panels a and c) and the corresponding values of $A_{V_{dir}}$ and $R_{V_{dir}}$ (panels b and d) for a simulated spiral galaxy (left plots) and a simulated elliptical galaxy (right plots). $A_{V_{dir}}$ and $R_{V_{dir}}$ were calculated from only the direct flux component, ignoring the scattered flux component. Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). An aperture radius of $20$ kpc was used. The galaxy dust density has been normalized so that the total optical depth along the R axis is $\tau_{V_{R}}=5$. The dashed black line reflects the true extinction $R_{V_{Zubko}}$.
@@
As was mentioned in Section 3.1, the trend between $A_{V_{eff}}$ and $R_{V_{eff}}$ is primarily driven by the scattering of light into the line of sight. This scattered component, when compared to the direct flux, is redder for the face-on observations, lowering the observed $R_{V_{eff}}$, and bluer for the edge-on view, resulting in higher $R_{V_{eff}}$. In addition to this, the fact that the sources now extend to different depths and contribute to the same optical path means that $R_{V_{eff}}$ deviates even more from $R_{V_{Zubko}}$.
@@
From the above, we conclude that the reddening caused by changes in the direct optical paths behaves similarly to the reddening caused by excess scattered light, with $R_{V_{eff}}$ increasing with the observation angle. In this case, this increase is due to the fact that, as we move from a face-on to an edge-on orientation, there is an increase in the number of different optical paths in our observation, as we are probing much deeper into the dust cloud. This also explains why the spiral galaxy face-on direct flux observations in Fig. 5 produce the closest results to $R_{V_{Zubko}}$, as this is the scenario in which the sources are most concentrated, given the narrower depth of the galaxy. We also find that, as is observed for the point source, the $A_{V_{eff}}$ values obtained for only the direct flux are systematically higher than those obtained from the full observed flux, as expected.
@@
This dust density effect is, however, still secondary to the variation in the observation angle, which produces a much larger change in $R_{V_{eff}}$. As such, the effective dust properties of a galaxy appear to be more a function of geometry and morphology than of the dust abundance itself. In practice, this means that, when using global galaxy aperture photometry, the true optical properties of the dust medium, namely $R_{V_{Zubko}}$, become obfuscated and cannot be easily recovered.
Figure 6: Values of $R_{V_{eff}}$ as a function of the total optical depth normalization along the R axis $\tau_{V_{R}}$, for a spiral (left) and elliptical galaxy (right). Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). An aperture radius of $20$ kpc was used. The dashed black line reflects the true extinction $R_{V_{Zubko}}$.
In Fig. 7, we present a comparison between the $A_{V_{eff}}$-$R_{V_{eff}}$ trend obtained from the various simulated observations detailed in Fig. 6 and the trend obtained for Dark Energy Survey (DES; Abbott et al., 2018) global galaxy fits, as discussed in Duarte et al. (2023). The plot covers different galaxy types, dust densities, and orientations. We find that, with the exception of the absolute $A_{V_{eff}}$ values, our simulations are able to fully reproduce the observed behavior. The mismatch in $A_{V}$ might actually stem from a bias in the DES $A_{V_{eff}}$ values of Duarte et al. (2023), which were obtained using the Prospector Python package (Johnson et al., 2021). These $A_{V_{eff}}$ values are found to be consistently higher than those recovered for similar galaxies by, for example, Zahid et al. (2013) and Salim et al. (2018), which are much closer to the values obtained from our simulations. We emphasize, however, that the relation between the two quantities is very well explained with our simulations. The increased level of scatter exhibited by the real data is most likely due to the fact that a real sample is expected to reflect a wide mixture of different morphologies and dust types.
Our results are also consistent with simulated observations of galaxies with different inclinations performed by Chevallard et al. (2013), as well as with other global galaxy observations (e.g, Salmon et al., 2016; Leja et al., 2017; Narayanan et al., 2018). Overall, we feel confident in asserting that the observed $A_{V_{eff}}$-$R_{V_{eff}}$ relation is indeed driven by the observation angle, with factors such as dust density and galaxy type contributing mainly to an increase in the trend dispersion.
@@ 3.2.2 Local observations
Contrary to what we found for a point source, observing with a $25$ pc resolution appears to be insufficient to meaningfully mitigate the scattered flux component across the entire galaxy, which now comes from integrated light. Even so, it appears that, for regions such as the centers of the galaxies, where the dust density is high, the scattered light is indeed negligible. In fact, looking at Fig. 9, where maps of $A_{V_{eff}}$ and $R_{V_{eff}}$ for these two galaxies are shown, we see that, for these central regions, we recover a value of $R_{V_{eff}}$ close to $R_{V_{Zubko}}$. As we move away from the center of the galaxy, we find a decrease in the values of $R_{V_{eff}}$. This is consistent with the gradient recovered by Hutton et al. (2015), who find that the observed $R_{V_{eff}}$ values decrease with the galactocentric distance, when using both ultraviolet and optical photometry. While the authors interpret this as evidence for a radial variation in intrinsic dust properties in galaxies, Fig. 9 shows that the same behavior can be recovered with a uniform dust type as a consequence of different levels of scattering and dust-to-stars distributions. There are also some outlier pixels with uncharacteristically high $R_{V_{eff}}$ values, but these arise mainly due to stochastic variations in the simulation and are particularly relevant for thin dust layers, such as the one present in the spiral edge-on view.
@@
Figure 9: Spatial maps of the values of $A_{V_{eff}}$ and $R_{V_{eff}}$, for a simulated spiral (left) and elliptical face-on galaxy (right). A pixel size corresponding to a $25$ pc resolution was used. The galaxy dust density has been normalized so that the total optical depth along the R axis is $\tau_{V_{R}}=5$. Only pixels with $A_{V_{eff}}>0$ and $R_{V_{eff}}>0$ are shown.
@@
Figure 10: 2D maps of the values for $A_{V_{eff}}$ and $R_{V_{eff}}$ for a face-on simulated spiral (left) and elliptical galaxy (right). The galaxy dust density has been normalized so that the total optical depth along the R axis is $\tau_{V_{R}}=5$. Each point corresponds to a single pixel with a resolution of $25$ pc, color-coded according to their galactocentric distance. Binned medians are shown in red, with error-bars defined by the bin standard deviation. The dashed black line reflects the true $R_{V}$ of the medium, with the dotted black lines defining a 10% deviation from this value. Only pixels with $A_{V_{eff}}>0$ and $R_{V_{eff}}>0$ are shown.
@@
The relation between the two dust parameters becomes even more apparent if we look at the $A_{V_{eff}}-R_{V_{eff}}$ trends for both face-on galaxy models, plotted in Fig. 10. As was stated above, in addition to the light scattered into the line of sight, which tends to be reddened and contributes to drive $R_{V_{eff}}$ down, one must also consider the existence of multiple sources at different depths in the same optical path, which, when taken together, contribute to an increase in $R_{V_{eff}}$. For higher values of $A_{V_{eff}}$, and thus higher dust densities, the amount of light that can escape from the innermost parts of the galaxy, be it scattered or otherwise, drastically goes down. As such, the observed flux becomes largely dominated by the direct flux of stars located in the outermost layers of the galaxy, which mitigates both of the above-mentioned effects.
@@
For this reason, we can recover a value of $R_{V_{eff}}$ within $10\%$ of $R_{V_{Zubko}}$ when observing areas with large optical depth near the center of the galaxies. This recovery is also aided by the high central source density, which also contributes to make the scattered component subdominant in this particular region. We end up slightly overestimating the true extinction value mainly because it is impossible to completely eliminate the contributions from sources deeper in the galaxy, which slightly increased the values of $R_{V_{eff}}$, as is seen for the stellar population modeled here.
@@
Figure 11: Similar to Fig. 10 for simulated edge-on spiral (left) and elliptical galaxy (right). For the spiral galaxy, the true extinction $R_{V}$ of the medium appears to be better traced by the low $A_{V_{eff}}$ regions, instead of the highly attenuated ones.